In an experiment using a linear air track a 0.5kg rider travelling to the right at 2m/s collides with a stationary rider of mass 1kg find the velocities of the two riders after collision if it is perfectly elastic


1
Expert's answer
2021-08-26T19:04:16-0400
"v_1=\frac{0.5-1}{0.5+1}2=-0.67\frac{m}{s}"

"v_2=\frac{2(0.5)}{0.5+1}2=1.33\frac{m}{s}"